# Skew brace extensions, second cohomology and complements

### v1

@@ Abstract.
We study extensions and second cohomology of skew left braces via the natural semi-direct products associated with the skew left braces. Let $0\to I\to E\to H\to 0$ be a skew brace extension and $\Lambda_{H}$ denote the natural semi-direct products associated with the skew left brace $H$. We establish a group homomorphism from $\operatorname{H}_{Sb}^{2}(H,I)$ into $\operatorname{H}_{Gp}^{2}(\Lambda_{H},I\times I)$, which turns out to be an embedding when $I\leq\operatorname{Soc}(E)$. In particular the Schur multiplier of a skew left braces $H$ embeds into the Schur multiplier of the group $\Lambda_{H}$. Analog of the Schur-Zassenhaus theorem is established for skew left braces in several specific cases. We introduce a concept called minimal extensions (which stay at the extreme end of split extensions) of skew left braces and derive many fundamental results. Several reduction results for split extensions of finite skew left braces by abelian groups (viewed as trivial left braces) are obtained.
@@ 1. Introduction
Cohomology of left braces and linear cycle sets with trivial action on an abelian group was introduced by Labed and Vendramin [19] . This was extended to the case when the action is non-trivial by Bardakov and Singh [7]. The concept of the (first and second) cohomology of skew left braces, which originated from the paper [1] by Bachiller, was formalised by the authors of the present article [25]. Connection with extension theory was developed. The aim of this article is to further advance this knowledge by exploiting the connection of skew left brace with the natural semi-direct product of the multiplicative group of the skew left brace by that of its additive group. Some reduction techniques for an exact sequence of skew braces by abelian $p$-groups are obtained.
@@
As a quick application of this theorem we get [21, Theorem 2.2] which states that the exponent of the Schur multiplier of a finite skew left brace $H$ divides $|H|^{2}$.
@@
Let $H$ be a skew left brace acting on a trivial brace $I$ via a good triplet of actions $(\nu,\mu,\sigma)$. Then $\operatorname{H}^{2}(H,I)\cong\operatorname{H}^{2}(H^{\operatorname{op}},I)$.
An ideal $I$ of a finite skew left brace $H$ is said to be a Hall ideal if the order of $I$ is coprime to the order of the quotient skew left brace $H/I$. We prove a brace analogous of the group theory result [4, 9.1.2] in several specific cases (Theorem 4.3), for example, when (i) $I\leq\operatorname{ker}(\lambda)$, (ii) $E$ is supersoluble, (iii) $(E,+)$ is a nilpotent group. We present various interesting results on Frattini sub-skew braces, minimal and split extensions of skew left braces. For example, we prove that the Frattini sub-skew brace of the direct sum of two finite skew left braces $E_{1}$ and $E_{2}$ is the direct sum of the Frattini sub-skew braces of $E_{1}$ and $E_{2}$.
In the last section we prove a reduction result for split skew brace extensions.
@@
Let $E$ be a brace admitting an ideal $I$ such that $I=I_{1}\times\cdots\times I_{r}$, where $I_{j}$ is an ideal of $E$ for each $1\leq j\leq r$. Let $I_{k}^{*}=\prod_{j\neq k}I_{j}$. Then the following hold true:
@@
We conclude this section by setting some notations. Note that additive and multiplicative identities of any skew brace coincide, which we denote by 0. A skew brace having cardinality $1$ is also mostly denoted by 0, instead of $\{0\}$. Inverse of an element $a\in A$, when considered in $(A,+)$ (respectively in $(A,\circ)$) is denoted by a-a (respectively by $a^{-1}$). The group commutator of two elements $a$ and $b$ of a skew brace $A$ in $(A,+)$ is denoted by $a,b]^{+}:=-a-b+a+b$, and in $(A,\circ)$, by $a,b]^{\circ}:=a^{-1}\circ b^{-1}\circ a\circ b$. The commutator $a,b]^{+}$ is also denoted by $\gamma_{+}(a,b)$. The commutators of larger weights are defined iteratively. We write $B\leq A$ to denote that $B$ is a sub-skew brace of the skew brace $A$, and $B\trianglelefteq A$ to denote that $B$ is an ideal of $A$. The quotient skew brace $A/I$ of $A$ by $A$ will be denoted by $\bar{A}$. The category of all skew left braces will be denoted by $\mathcal{SB}$.
@@ 2. Preliminaries
    If $E$ is an extension of $H$ by a trivial brace $I$, then the extension $\Lambda_{E}$ is an abelian extension of $\Lambda_{H}$ by $I\times I$. Furthermore, if $I\leq\operatorname{Ann}(E)$, then the induced extension of $\Lambda$-groups is a central extension as $\Lambda_{\operatorname{Ann}(E)}\leq\operatorname{Z}(\Lambda_{E})$ (see [24, Proposition 3.4] for details).
@@ 3. From skew braces to groups : extensions and cohomology
Let $\operatorname{Z}^{2}_{Sb}(H,I)$ denote the set of all brace $2$-cocycles with coefficients in $I$, and $\operatorname{Z}^{2}_{Gp}((H,+),I)$ and $\operatorname{Z}^{2}_{Gp}((H,\circ),I)$ denote the set of group $2$-cocycles of the groups $(H,+)$ and $(H,\circ)$ respectively with coefficients in $I$. Then it follows from [25, Page 10, Eqn. (13)] that $(\beta,\tau)\in\operatorname{Z}^{2}_{Sb}(H,I)$ if and only if $\beta\in\operatorname{Z}^{2}_{Gp}((H,+),I)$ and $\overline{\tau}\in\operatorname{Z}^{2}_{Gp}((H,\circ),I)$ and the following equation holds:
@@
To defend our statement about the usefulness of the preceding result, we derive [21, Theorem 2.2].
@@
The preceding result also follows from the following much stronger result proved in [21, Proposition 2.4].
@@
Let $(H,+,\circ)$ be a brace of order $p^{n}$, with $p$ an odd prime. It follows from [11, Section 2] that
@@
We can define a homomorphism from $\operatorname{H}^{2}_{Sb}(H,I)$ to $\operatorname{H}^{2}_{Gp}((H,\circ),I)$; $(\beta,\tau)]\mapsto[\overline{\tau}]$. It is not difficult to show that $\bar{\tau}=\nu^{-1}_{h_{1}\circ h_{2}}\tau$ is a group $2$-cocycle of $(H,\circ)$ with coefficients in $I$ and action $\sigma$. If $(\beta,\tau)]\mapsto[\overline{\tau}]$ is well defined, then, being the projection map, it will a group homomorphism. So we only need to show that the map is well defined, that is, if $(\beta_{1},\tau_{1})$ and $(\beta_{2},\tau_{2})$ are cohomologous, then so are the group $2$-cocycles $\bar{\tau}_{1}$ and $\bar{\tau}_{2}$. Assume that $(\beta_{1},\tau_{1})$ and $(\beta_{2},\tau_{2})$ are cohomologous. Then there exists a map $\theta:H\to I$ such that (see [25, proof of Theorem 3.6])
@@
It then follows from [22, Proposition 2.2] that the second cohomology groups
$\operatorname{H}^{2}\big(\operatorname{Diag}(\Lambda_{H}),J\big)\cong\operatorname{H}^{2}\big((H,\circ),I\big)$
@@
Putting all together we get the following result, which generalizes [21, Theorem 1.37].
@@
    The pair $(\beta,\tau)\in\operatorname{Z}^{2}(H,I)$ if and only if $(\beta^{\operatorname{op}},\tau^{\operatorname{op}})\in\operatorname{Z}^{2}(H^{\operatorname{op}},I)$.
@@
    The map $\tilde{\varphi}:\operatorname{H}^{2}(H,I)\rightarrow\operatorname{H}^{2}(H^{\operatorname{op}},I)$ given by $(\beta,\tau)]\mapsto[(\beta^{\operatorname{op}},\tau^{\operatorname{op}})]$ is an isomorphism.
@@ 4. Complements and supplements
A sub-skew brace $K$ of a skew brace $E$ is said to be a Hall $\pi$-sub-skew brace if $K$ is a $\pi$-number and the index of $K$ in $E$ is a $\pi^{\prime}$-number. If $\pi=\{p\}$, then Hall $\pi$-sub-skew brace $K$ of $E$ is said to be Sylow $p$-sub-skew brace of $E$. We first state the following result [10, Theorem 1.6]:
@@
Let $H$ be a supersoluble skew brace and let $I$ be an abelian group such that $|H|$ and $|I|$ are coprime. Then $\operatorname{H}^{2}_{Sb}(H,I)$ is trivial. In particular, if $H$ has squarefree order and $\gcd(|H|,|I|)=1$, then $\operatorname{H}^{2}_{Sb}(H,I)$ is trivial.
@@
Proposition 4.5.
Let $\operatorname{H}^{2}_{Sb}(H,I)$ be the second cohomology group of $H$ by $I$, where $H$ and $I$ are skew braces of coprime orders with associated action $(\mu,\nu,\sigma)$. Then $\operatorname{H}^{2}_{Sb}(H,I)$ is trivial if any of the following holds:
@@
If $\mu$ is trivial, then every element of $\operatorname{H}^{2}_{Sb}(H,I)$ corresponds to an extension $E$ such that $I\leq\operatorname{Z}(E,+)$. It is well known that a Hall normal subgroup contained in the center admits a unique complement; hence, the complement of $I$ in $(E,+)$ is characteristic and therefore forms a skew brace.
The second assertion follows from Theorem 4.3(3). ∎
@@
Next, we state a characterization of the minimal ideal of a finite soluble skew brace, established in [4, Theorem B], which we’ll use in the next result.
@@
Proof.
Let the order $E$ be $p_{1}^{k_{1}}\cdots p_{r}^{k_{r}}$, where $p_{i}$’s are distinct primes and $k_{i}$’s and $r$ are natural numbers. We proceed by the method of contradiction. So assume that $E$ is the skew brace of the smallest order not admitting a sub-skew brace of order $p_{j}^{k_{j}}$ for at least one $j$, but all skew braces $A$ of order smaller then the order of $E$ admits Sylow $p$-sub-skew brace for every prime $p$ dividing the order of $A$. By Theorem 4.7, $E$ admits a minimal ideal $N$ of prime power order $p^{k}$ for a prime $p$ and a positive integer $k\geq 1$. So $p=p_{i}$ for some $i$. Consider the quotient skew brace $E/N$. Since $|E/N|<|E|$ and the hypothesis of the theorem holds true for $E/N$, the assertion holds true for $E/N$.
Let $\bar{K_{j}}$ denote a sub-skew brace of $E/N$ having order $p_{j}^{k_{j}}$ for $j\neq i$. Note that, for the prime $p_{i}$, the sub-skew brace $K_{i}$ of $E/N$ has order $|\bar{K_{i}}|=p_{i}^{k_{i}-k}$. If $k_{i}=k$, then $N$ itself is the Sylow $p_{i}$-sub-skew brace of $E$ and we get an exact sequence $0\to N\to E\to E/N\to 0$. Since $|N|$ is coprime to $|E/N|$, by Theorem 4.3(3), $N$ admits a complement $L$ in $E$. Since $L<|E|$ and also the given hypothesis holds true for $L$, it admits Sylow $p_{j}$-sub-skew brace of $E$ for all $j\neq i$. Now assume that $k<k_{i}$. Let $K_{j}$ be the pre-image of $\bar{K_{j}}$, $1\leq j\leq r$, in $E$. Then, obviously, $K_{i}$ is a Sylow $p_{i}$-sub-skew brace of $E$. For $j\neq i$, we get the exact sequence $0\to N\to K_{j}\to\bar{K_{j}}\to 0$. Since the order of $N$ is coprime to the order of $\bar{K_{j}}$, by Theorem 4.3(3), $N$ admits a complement $L_{j}$ in $K_{j}$. It is now clear that $L_{j}$ is a Sylow $p_{j}$-sub-skew brace of $E$ for all $j\neq i$. Hence $E$ admits a Sylow $p$-sub-skew brace for every prime divisor $p$ of $|E|$, which contradicts the minimality of $E$. The proof is complete. ∎
Next, we examine the structure of supersoluble skew braces of square-free order. As mentioned above, the semidirect product of a skew brace $H$ by a skew brace $I$ will be denoted by $I\rtimes_{Sb}H$. First, we recall a basic result from [5, Corollary 3.9].
@@
Next, we recall a description of split extensions of skew braces from [23, Theorem 3.1, Theorem 3.3].
@@
Let $I$ be an abelian sub-skew brace of $E$ which admits a supplement $G$. Then $J:=I\cap G$ is an ideal of $E$ ([4, Lemma 28]). Now $G/J$ is a complement of $I/J$ in $E/J$.
@@
as $\Phi(H/(H\cap I))$ is the intersection of all maximal sub-skew braces of $H$ containing $H\cap I$. But using part (i), $H\cap I$ is already contained in all maximal sub-skew braces of $H$. Hence
@@
The following are two more easy consequences of Proposition 4.16 on simple extensions.
@@
Let $E$ be a minimal extension of a skew brace $H$ with respect to $I$ and let $J\leq I$ be such that $J$ is an ideal of $E$. Then $E/J$ is minimal extension of $H$ with respect to $I/J$.
@@
Suppose that $E$ is not a minimal extension. Then $I$ admits a proper supplement $G$, say in $E$. Since $I$ is an ideal of $E$, $G\cap E$ is an ideal of $G$. We claim that $G\cap I$ is an ideal of $E$. Let any $x\in G\cap I$, and $h\in E$, then we know that $h=g\circ z$ for some $g\in G$, $z\in I$. Then
@@
Let $x\in K_{n}$. Then there exist elements $y_{1},y_{2}\in K_{n-1}$ such that $x=\epsilon\epsilon^{\prime}(y_{1},y_{2})$, where $\epsilon\in\{+,\bar{+},\circ,\bar{\circ}\}$, $\epsilon^{\prime}\in\{1,\delta_{1},\delta_{2},\chi_{1},\chi_{2}\}$ and $1_{M}$ denotes the identity element of $M$.
@@
    For every $1\neq x\in\mathcal{W}(X)$, there exist an integer $n\geq 0$, elements ϵ_0, ϵ_1, …, ϵ_(n-1) ∈{+, ¯+, , ¯∘} and ϵ’_0, ϵ’_1, …, ϵ’_(n-1) ∈{1_M, δ_1, δ_2, χ_1, χ_2} such that x = ϵ_(n-1) ϵ’_(n-1) ((ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (, ), ϵ_0 ϵ’_0 (, ))), (ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (, ), ϵ_0 ϵ’_0 (, )))), where $\bullet\in X\cup\{0\}$ and may take different values at different positions. Also, for any brace homomorphism $\phi:A\to B$, the image of $x$ under $\phi$, denoted by $\phi(x)$, is given by ϵ_(n-1) ϵ’_(n-1) ((ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (’, ’), ϵ_0 ϵ’_0 (’, ’))), (ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (’, ’), ϵ_0 ϵ’_0 (’, ’)))), where $\bullet^{\prime}=\phi(\bullet)\in B$.
@@
The second assertion follows from the definition of $\mathcal{W}(X)$, definition of a brace homomorphism and applying an easy induction. Indeed, for the induction base, we see that $x=+(0,x)$, $x=+\delta_{2}(0,x)$ and $x^{-1}=\circ\chi_{2}(0,x)$ for the elements of $X_{1}$. Note that these expressions are not unique. So the base case when $1\neq x\in K_{1}$ is clear. ∎
We remark that the element $1\neq x\in\mathcal{W}(X)$ as in the preceding proposition can be expressed as
@@
where $a_{n-1}\in K_{n-1},a_{n-2}\in K_{n-2},\ldots,a_{1}\in K_{1},a_{01},a_{02}\in X$. For the convenience of the notation, let us denote this expression in word map form as follows: Let $1\neq x\in\mathcal{W}(X)$ as in the preceding proposition. Then we can write $x$ as an image of a word $w:X^{m}\to\mathcal{W}(X)$, that is, $x=w(x_{i_{1}},\ldots,x_{i_{m}})$, where $m\leq 2^{n}$ and $x_{i_{j}}\in X$. More precisely, $w$ is the expression
@@
where $\bullet$’s are indeterminant $t_{1},t_{2},\ldots,t_{m}$.
@@
Now let $a\in E$ be arbitrary. We know that $a=s(x)+y$ for some $x\in H$ and $y\in I$. Writing $x=w(x_{i_{1}},\ldots,x_{i_{m}})$ and using (4.4), we get

### v2

@@ Abstract.
We study extensions and the second cohomology groups of skew left braces via the associated natural semi-direct product groups. Let $0\to I\to E\to H\to 0$ be a skew brace extension and $\Lambda_{H}$ denote the natural semi-direct product associated with the skew left brace $H$. We establish a group homomorphism from $\operatorname{H}_{Sb}^{2}(H,I)$, the second cohomology group of braces, into $\operatorname{H}_{Gp}^{2}(\Lambda_{H},I\times I)$, the second cohomology group of groups, which turns out to be an embedding when $I\leq\operatorname{Soc}(E)$, the socle of $E$. In particular, the Schur multiplier of a skew left brace $H$ embeds into the Schur multiplier of the group $\Lambda_{H}$. Analog of the Schur-Zassenhaus theorem is established for skew left braces in several specific cases. We introduce a concept called minimal extensions (which stay at the extreme end of split extensions) of skew left braces and derive many fundamental results. Several reduction results for split extensions of finite skew left braces by abelian groups (viewed as trivial left braces) are obtained.
@@ 1. Introduction
Cohomology of left braces and linear cycle sets with trivial action on an abelian group was introduced by Labed and Vendramin [15] . This was extended to the case when the action is non-trivial by Bardakov and Singh [6]. The concept of the (first and second) cohomology of skew left braces, originated from the article [1] by Bachiller, was formalised by the authors of the present article in [20]. Connection with extension theory was developed. The aim of this article is to further advance this knowledge by exploiting the connection of skew left brace with the natural semi-direct product of the multiplicative group of the skew left brace by that of its additive group. Several foundational results on minimal extensions of skew left braces and Frattini sub-skew braces are produced. Some reduction techniques for an exact sequence of skew left braces by abelian $p$-groups to split are obtained.
@@
As a quick application of this theorem we get [16, Theorem 2.2] which states that the exponent of the Schur multiplier of a finite skew left brace $H$ divides $|H|^{2}$.
@@
Let $H$ be a skew left brace acting on a trivial brace $I$ via a good triplet of actions $(\nu,\mu,\sigma)$. Then $\operatorname{H}^{2}_{Sb}(H,I)\cong\operatorname{H}^{2}_{Sb}(H^{\operatorname{op}},I)$.
An ideal $I$ of a finite skew left brace $H$ is said to be a Hall ideal if the order of $I$ is coprime to the order of the quotient skew left brace $H/I$. We prove a brace analogous of the group theory result [3, 9.1.2] in several specific cases (Theorem 4.3), for example, when (i) $I\leq\operatorname{ker}(\lambda)$, (ii) $E$ is supersoluble, (iii) $(E,+)$ is a nilpotent group. We present various interesting results on Frattini sub-skew braces, minimal and split extensions of skew left braces. For example, we prove that the Frattini sub-skew brace of the direct sum of two finite skew left braces $E_{1}$ and $E_{2}$ is the direct sum of the Frattini sub-skew braces of $E_{1}$ and $E_{2}$.
In the last section we prove a reduction result for split brace extensions.
@@
Let $E$ be a left brace admitting an ideal $I$ such that $I=I_{1}\times\cdots\times I_{r}$, where $I_{j}$ is an ideal of $E$ for each $1\leq j\leq r$. Let $I_{k}^{*}=\prod_{j\neq k}I_{j}$. Then the following hold true:
@@
We conclude this section by setting some notations. Note that additive and multiplicative identities of any skew left brace coincide, which we denote by 0. A skew left brace having cardinality $1$ is also mostly denoted by 0, instead of $\{0\}$. Inverse of an element $a\in A$, when considered in $(A,+)$ (respectively in $(A,\circ)$) is denoted by a-a (respectively by $a^{-1}$). The group commutator of two elements $a$ and $b$ of a skew left brace $A$ in $(A,+)$ is denoted by $a,b]^{+}:=-a-b+a+b$, and in $(A,\circ)$, by $a,b]^{\circ}:=a^{-1}\circ b^{-1}\circ a\circ b$. The commutator $a,b]^{+}$ is also denoted by $\gamma_{+}(a,b)$. The commutators of larger weights are defined iteratively. We write $B\leq A$ to denote that $B$ is a sub-skew brace of the skew brace $A$, and $B\trianglelefteq A$ to denote that $B$ is an ideal of $A$. The quotient skew brace $A/I$ of $A$ by $A$ will be denoted by $\bar{A}$. Isomorphism theorems for skew braces were established in [5], which we’ll use without further reference. The category of all skew left braces will be denoted by $\mathcal{SB}$.
@@ 2. Preliminaries
    If $E$ is an extension of $H$ by a trivial brace $I$, then the extension $\Lambda_{E}$ is an abelian extension of $\Lambda_{H}$ by $I\times I$. Furthermore, if $I\leq\operatorname{Ann}(E)$, then the induced extension of $\Lambda$-groups is a central extension as $\Lambda_{\operatorname{Ann}(E)}\leq\operatorname{Z}(\Lambda_{E})$ (see [19, Proposition 3.4] for details).
@@ 3. From skew braces to groups : extensions and cohomology
Let $\operatorname{Z}^{2}_{Sb}(H,I)$ denote the set of all brace $2$-cocycles with coefficients in $I$, and $\operatorname{Z}^{2}_{Gp}((H,+),I)$ and $\operatorname{Z}^{2}_{Gp}((H,\circ),I)$ denote the set of group $2$-cocycles of the groups $(H,+)$ and $(H,\circ)$ respectively with coefficients in $I$. Then it follows from [20, Page 10, Eqn. (13)] that $(\beta,\tau)\in\operatorname{Z}^{2}_{Sb}(H,I)$ if and only if $\beta\in\operatorname{Z}^{2}_{Gp}((H,+),I)$ and $\overline{\tau}\in\operatorname{Z}^{2}_{Gp}((H,\circ),I)$ and the following equation holds:
@@
To defend our statement about the usefulness of the preceding result, we derive [16, Theorem 2.2].
@@
The preceding result also follows from the following much stronger result proved in [16, Proposition 2.4].
@@
Let $(H,+,\circ)$ be a brace of order $p^{n}$, with $p$ an odd prime. It follows from [10, Section 2] that
@@
We can define a homomorphism from $\operatorname{H}^{2}_{Sb}(H,I)$ to $\operatorname{H}^{2}_{Gp}((H,\circ),I)$; $(\beta,\tau)]\mapsto[\overline{\tau}]$. It is not difficult to show that $\bar{\tau}=\nu^{-1}_{h_{1}\circ h_{2}}\tau$ is a group $2$-cocycle of $(H,\circ)$ with coefficients in $I$ and action $\sigma$. If $(\beta,\tau)]\mapsto[\overline{\tau}]$ is well defined, then, being the projection map, it will a group homomorphism. So we only need to show that the map is well defined, that is, if $(\beta_{1},\tau_{1})$ and $(\beta_{2},\tau_{2})$ are cohomologous, then so are the group $2$-cocycles $\bar{\tau}_{1}$ and $\bar{\tau}_{2}$. Assume that $(\beta_{1},\tau_{1})$ and $(\beta_{2},\tau_{2})$ are cohomologous. Then there exists a map $\theta:H\to I$ such that (see [20, proof of Theorem 3.6])
@@
It then follows from [17, Proposition 2.2] that the second cohomology groups
$\operatorname{H}^{2}_{Gp}\big(\operatorname{Diag}(\Lambda_{H}),J\big)\cong\operatorname{H}^{2}_{Gp}\big((H,\circ),I\big)$
@@
Putting all together we get the following result, which generalizes [16, Theorem 1.37].
@@
    The pair $(\beta,\tau)\in\operatorname{Z}^{2}_{Sb}(H,I)$ if and only if $(\beta^{\operatorname{op}},\tau^{\operatorname{op}})\in\operatorname{Z}^{2}_{Sb}(H^{\operatorname{op}},I)$.
@@
    The map $\tilde{\varphi}:\operatorname{H}^{2}_{Sb}(H,I)\rightarrow\operatorname{H}^{2}_{Sb}(H^{\operatorname{op}},I)$ given by $(\beta,\tau)]\mapsto[(\beta^{\operatorname{op}},\tau^{\operatorname{op}})]$ is an isomorphism.
@@ 4. Complements and supplements
A sub-skew brace $K$ of a skew brace $E$ is said to be a Hall $\pi$-sub-skew brace if $K$ is a $\pi$-number and the index of $K$ in $E$ is a $\pi^{\prime}$-number. If $\pi=\{p\}$, then Hall $\pi$-sub-skew brace $K$ of $E$ is said to be Sylow $p$-sub-skew brace of $E$. We first state the following result [9, Theorem 1.6]:
@@
Let $H$ be a finite supersoluble skew brace and let $I$ be a finite abelian group such that $|H|$ and $|I|$ are coprime. Then $\operatorname{H}^{2}_{Sb}(H,I)$ is trivial. In particular, if $H$ has squarefree order and $\gcd(|H|,|I|)=1$, then $\operatorname{H}^{2}_{Sb}(H,I)$ is trivial.
@@
Corollary 4.5.
Let $H$ be a finite skew brace and let $I$ be a finite abelian group such that $|H|$ and $|I|$ are coprime. Then $\operatorname{H}^{2}_{Sb}(H,I)$, with associated action $(\mu,\nu,\sigma)$, is trivial if any of the following holds:
@@
If $\mu$ is trivial, then every element of $\operatorname{H}^{2}_{Sb}(H,I)$ corresponds to an extension $E$ of $H$ by $I$ such that $I\leq\operatorname{Z}(E,+)$. It is well known that a Hall normal subgroup contained in the center admits a unique complement, which by Theorem 4.3(7) is a left ideal of $E$. This proves that $I$ admits a complement in $E$, and therefore $\operatorname{H}^{2}_{Sb}(H,I)$ is trivial.
Let $E$ be an extension of $H$ by $I$. Assume that $\mu_{h}(x)=\nu_{h}(\sigma_{h}(x))$ for all $h\in H$ and $x\in I$. Using this equation in (3), we immediately get $I\leq\operatorname{ker}(\lambda)$ in $E$. The assertion now follows from Theorem 4.3(3). ∎
@@
Next, we state a characterization of the minimal ideal of a finite soluble skew brace, established in [3, Theorem B], which we’ll use in the next result.
@@
Let $E$ be a finite soluble skew brace such that $\lambda_{x}(y)$ commutes with $y$ for all $x,y\in E$. Then $E$ admits a Hall $\pi$-sub-skew brace, for every subset $\pi$ of the set of all prime divisors of $|E|$.
Proof.
Let the order of $E$ be $p_{1}^{k_{1}}\cdots p_{r}^{k_{r}}$, where $p_{i}$’s are distinct primes and $k_{i}$’s and $r$ are natural numbers. Note that every sub-skew brace and every quotient brace of $E$ satisfy the hypothesis of the theorem. We proceed by the method of contradiction. So assume that $E$ is a skew brace of the smallest order satisfying the given hypothesis, but not admitting a Hall $\pi^{\prime}$-sub-skew brace for at least one non-empty subset $\pi^{\prime}$ of the set $\{p_{1},\ldots,p_{r}\}$. By Theorem 4.7, $E$ admits a minimal ideal $N$ of prime power order $p^{k}$ for a prime $p$ and a positive integer $k\geq 1$. So $p=p_{i}$ for some $i$. Consider the quotient skew brace $E/N$. Since $|E/N|<E|$ and the hypothesis of the theorem holds true for $E/N$, the assertion holds true for $E/N$.
Let $\bar{K_{\pi}}$ denote some Hall $\pi$-sub-skew brace of $E/N$ for each $\pi$. Let $\{p_{i}\}\subset\pi^{\prime\prime}$. If $\pi^{\prime\prime}=\{p_{i}\}$ and $N$ itself is a Hall $\pi^{\prime\prime}$-sub-skew brace of $E$, then we get an exact sequence $0\to N\to E\to E/N\to 0$. Since $|N|$ is coprime to $|E/N|$, by Theorem 4.3(3), $N$ admits a complement $L$ in $E$. Since $L$ is a proper sub-skew brace $E$, it admits Hall $\pi$-sub-skew brace, which is also Hall $\pi$-sub-skew brace of $E$, for all $\pi$. Now assume that $N$ is not a Hall $\pi^{\prime\prime}$-sub-skew brace of $E$. Let $K_{\pi}$ be the pre-image of $\bar{K_{\pi}}$, for each $\pi$, in $E$. Then, obviously, $K_{\pi^{\prime\prime}}$ is a Hall $\pi^{\prime\prime}$-sub-skew brace of $E$. For $\pi\neq\pi^{\prime\prime}$, we get the exact sequence $0\to N\to K_{\pi}\to\bar{K_{\pi}}\to 0$. Since the order of $N$ is coprime to the order of $\bar{K_{\pi}}$, by Theorem 4.3(3), $N$ admits a complement $L_{\pi}$ in $K_{\pi}$. It is now clear that $L_{\pi}$ is a Hall $\pi$-sub-skew brace of $E$ for all $\pi\neq\pi^{\prime\prime}$. This contradicts the minimality of $E$, and the proof is complete. ∎
As a special case of the preceding theorem, we immediately get
Corollary 4.9.
@@
Next, we examine the structure of supersoluble skew braces of square-free order. As mentioned above, the semidirect product of a skew brace $H$ by a skew brace $I$ will be denoted by $I\rtimes_{Sb}H$. First, we recall a basic result from [4, Corollary 3.9].
@@
Next, we recall a description of split extensions of skew braces from [18, Theorem 3.1, Theorem 3.3].
@@
Let $I$ be an abelian sub-skew brace of $E$ which admits a supplement $G$. Then $J:=I\cap G$ is an ideal of $E$ ([3, Lemma 28]). Now $G/J$ is a complement of $I/J$ in $E/J$.
@@
where $\Phi(H/(H\cap I))$ is the intersection of all maximal sub-skew braces of $H$ containing $H\cap I$. But using part (i), $H\cap I$ is already contained in all maximal sub-skew braces of $H$. Hence
@@
The following two results are other easy consequences of Proposition 4.17 on simple extensions.
@@
Let $E$ be a minimal extension of a skew brace $H$ with respect to $I$ and let $J\leq I$ be such that $J$ is an ideal of $E$. Then $E/J$ is a minimal extension of $H$ with respect to $I/J$.
@@
Suppose that $E$ is not a minimal extension. Then $I$ admits a proper supplement $G$, say, in $E$. Since $I$ is an ideal of $E$, $G\cap I$ is an ideal of $G$. We claim that $G\cap I$ is an ideal of $E$. For any $h\in E$, we know that $h=g\circ z$ for some $g\in G$, $z\in I$. Then, for any $x\in G\cap I$, we get
@@
Let $x\in K_{n}$. Then there exist elements $y_{1},y_{2}\in K_{n-1}$ such that $x=\epsilon\epsilon^{\prime}(y_{1},y_{2})$, where $\epsilon\in\{+,\bar{+},\circ,\bar{\circ}\}$, $\epsilon^{\prime}\in\{1_{M},\delta_{1},\delta_{2},\chi_{1},\chi_{2}\}$, $1_{M}$ being the identity element of $M$.
@@
    For every $0\neq x\in\mathcal{W}(X)$, there exist an integer $n\geq 0$, elements ϵ_0, ϵ_1, …, ϵ_(n-1) ∈{+, ¯+, , ¯∘} and ϵ’_0, ϵ’_1, …, ϵ’_(n-1) ∈{1_M, δ_1, δ_2, χ_1, χ_2} such that x = ϵ_(n-1) ϵ’_(n-1) ((ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (, ), ϵ_0 ϵ’_0 (, ))), (ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (, ), ϵ_0 ϵ’_0 (, )))), where $\bullet\in X\cup\{0\}$ and may take different values at different positions. Also, for any brace homomorphism $\phi:A\to B$, the image of $x$ under $\phi$, denoted by $\phi(x)$, is given by ϵ_(n-1) ϵ’_(n-1) ((ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (’, ’), ϵ_0 ϵ’_0 (’, ’))), (ϵ_1 ϵ’_1 (ϵ_0 ϵ’_0 (’, ’), ϵ_0 ϵ’_0 (’, ’)))), where $\bullet^{\prime}=\phi(\bullet)\in B$.
@@
The second assertion follows from the definition of $\mathcal{W}(X)$, definition of a brace homomorphism and applying an easy induction. Indeed, for the induction base, we see that $x=+(0,x)$, $x=+\delta_{2}(0,x)$ and $x^{-1}=\circ\chi_{2}(0,x)$ for the elements of $X_{1}$. Note that these expressions are not unique. So the base case when $0\neq x\in K_{1}$ is clear. ∎
We remark that the element $0\neq x\in\mathcal{W}(X)$ as in the preceding proposition can be expressed as
@@
where $a_{n-1}\in K_{n-1},a_{n-2}\in K_{n-2},\ldots,a_{1}\in K_{1},a_{01},a_{02}\in X$. For the convenience of the notation, let us denote this expression in word map form as follows: Let $0\neq x\in\mathcal{W}(X)$ as in the preceding proposition. Then we can write $x$ as an image of a word $w:X^{m}\to\mathcal{W}(X)$, that is, $x=w(x_{i_{1}},\ldots,x_{i_{m}})$, where $m\leq 2^{n}$ and $x_{i_{j}}\in X$. More precisely, $w$ is the expression
@@
where $\bullet$’s are indeterminants $t_{1},t_{2},\ldots,t_{m}$.
@@
Now let $a\in E$ be an arbitrary element. We know that $a=s(x)+y$ for some $x\in H$ and $y\in I$. Writing $x=w(x_{i_{1}},\ldots,x_{i_{m}})$ and using (4.4), we get